FERMI RESONANCE PERTURBATIONS OF ORBITAL ANGULAR MOMENTUM (RENNER-TELLER) COUPLING IN TRIATOMIC MOLECULES: THE STRETCH-BENDER APPROACH

\underline{G. DUXBURY}^{a}, B. D. MCDONALD, Department of Physics and Applied Physics, Strathclyde University, Glasgow G4 0NG, Scotland, U.K.; Ch. JUNGEN, Laboratoire Aimée Coton de CNRS, Université de Paris-Sud, 91405 Orsay CEDEX, FRANCE; A. ALIJAH, Facultät für Chemie, Universität Bielefeld, D-4800 Bielefeld, Germany.

Fermi resonance effects on Renner-Teller coupling in NH_{2} and CH_{2} have been identified. We have developed an extension of the semi-rigid bender approach of Jungen and Merer^{b} and of Duxbury and Dixon^{c} to calculate the interplay between Renner-Teller and vibrational resonances in detail. The basic idea is that if in the semi-rigid bender the bond length relaxes to follow the minimum on the potential energy surface, the effect of the stretchbend coupling should be minimised in this basis. This should simplify the calculations compared to those using a full three dimensional basis when only a few quanta of the stretching vibrations are excited via resonances. The effects of the interplay between Renner-Teller and vibrational coupling will be discussed using some of the most recent data^{d} obtained for the b̃ state of CH_{2}, making use of two-dimensional wavefunctions without and with Fermi resonance effects.

avisiting fellow at JILA

bMolec. Phys. 40, 1 (1980)

cMolec. Phys. 43, 255 (1981)

dB. C. Chang, M. Wu, G. E. Hall and T. J. Sears J. Chem. Phys. 101, 9236 (1994)